OpenStudy (anonymous):

find a polynomial function with integer coefficients that has the given zeros. 2, 4+i, 4-i

OpenStudy (anonymous):

If a, b and c are the zeroes of the polynomial, then the polynomial is given by: \[(x-a)(x-b)(x-c) = 0\] Use this concept to find the polynomial..

OpenStudy (shubhamsrg):

since x=2,x=4+i and x=(4-i) satisfy the req polynomial, then (x-2)(x- (4+i)) (x- (4-i)) =0 just expand LHS alternative can be assume the eqn to be x^3 + px^2 + qx + r =0 then sum of roots = -p product of roots in pairs of 2 = q product of roots = -r try now..

OpenStudy (anonymous):

i set it up already. i need help multiplying it together :p

OpenStudy (shubhamsrg):

wolfram alpha might help you then..hmm